OpenStudy (iholman):

A triangular pyramid has the same base and height as a triangular prism. The volume of the pyramid is 26 cubic centimeters. What is the volume of the triangular prism? cm3

Directrix (directrix):

The volume of the prism is the area of its base times its height. A = B*h The volume of the pyramid is 1/3 * B *h. These two have the same base area and height in your problem.

Directrix (directrix):

V = B*h V = 1/3 * B *h The pyramid's volume is 1/3 that of the prism.

Directrix (directrix):

Volume of the pyramid is 26 cubic centimeters. So, what is the volume of the prism? @iholman

Directrix (directrix):

What is 26 * 3 =

OpenStudy (iholman):

78

OpenStudy (iholman):

78

Directrix (directrix):

Correct. 78 cm^3

OpenStudy (iholman):

What is the lateral surface area of the triangular pyramid? Use the net of the pyramid to find the lateral surface area. cm2 Triangular pyramid with equilateral triangle base and its net - middle, left, right triangles base 2 feet, top triangle height 5 feet